OpenStudy (anonymous):

find the following product: -3z(z^3-3z^2+6) I am having such a hard time with this one can someone figure this out???? PLEASE

OpenStudy (anonymous):

Think distributive property; go one term at a time.

OpenStudy (anonymous):

Can you please show me?

OpenStudy (anonymous):

Take the -3z and multiply it by each of the terms in the parentheses: -3z(z^3-3z^2+6) = -3z*z^3 + -3z*-3z^2 + -3z*6

OpenStudy (anonymous):

Then how would you break it down? I am not good at this at all... I am sorry.

OpenStudy (anonymous):

just multiply the stuff together: -3z*z^3 = -3z^4. I'm sure you can figure out the rest.

OpenStudy (anonymous):

I am assuming it would be -3z(z^3-3z^2+6)=-3z^4+9z^3-18z?